find central angle theta which forms a sector of area 9 square inches of a circle of radius 5 inches. answer is being asked in radians. Please help.
Answer by lwsshak3(11628)

find central angle theta which forms a sector of area 9 square inches of a circle of radius 5 inches. answer is being asked in radians.
***
Area of sector=(1/2)r^2*theta=9
theta=9/(r^2/2)=9/(25/2=18/25
theta=18/25 radians